# B. Nikolić

### v1

@@ 5 Concluding remarks
In this paper the pure spinor model of type II superstring theory is studied. We choose to work with the superstring propagating in the presence of the constant background fields. The context of studying is T-dualization - combined bosonic and fermionic T-duality and double space representation of the full (bosonic++fermionic) T-dualization.
@@
At the beginning there are the separate recapitualations of the known resulats for bosonic and fermionic T-dualization of the type II superstring theory with constant background fields. The short discussion about choice of background fields is given and it is shown briefly the accordance with the consistency conditions. The main steps and results from [11] are repeated. Also in these papers [11] one can find more detailed analysis regarding the choice of the background fields.
@@
In the next part of the paper we showed that bosonic and fermionic Buscher T-dualization procedure, applied on the type II superstring with constant background fields, commute. Independent on sequence of the application of T-duality procedures, the result is unique - full T-dualized theory. Further, the background fields of the full T-dualized theory are obtained. The result is not trivial in the sense that we work with the model under approximations - all background fields are constant. This result is an additional proof that such choice is coorrect.
@@
Final part of the paper is dedicated to the representation of full T-dualization in double space. Here we introduced the double coordinate and define the matrix of the full T-dualization, $\mathcal{T}^{M}{}_{N}$, which exchanges $x^{\mu}\leftrightarrow y_{\mu}$, $\theta^{\alpha}\leftrightarrow z_{\alpha}$ and $\bar{\theta}^{\alpha}\leftrightarrow\bar{z}_{\alpha}$, simultaneously. The full T-dualized coordinate is defined as $^{\star}Z^{M}=\mathcal{T}^{M}{}_{N}Z^{N}$. The next step was to rewrite the T-dual transformation laws using double coordinate and introducing so called generalized metric $\mathcal{H}_{\pm MN}$. Note that every chiral sector has its own generalized metric. Demanding that initial double coordinate $Z^{M}$ and fully T-dualized $^{\star}Z^{M}$ have the same T-dual transformation laws we get the form of the T-dualized generalized metric, $^{\star}\mathcal{H}_{\pm}=\mathcal{T}^{T}\mathcal{H}_{\pm}\mathcal{T}$. From one of these two equations we calculate the background fields of the fully T-dualized theory which are in full accordance with those obtained in the analytical (Buscher) procedure. For this specific choice of background fields we proved the equivalence of the analytical and double space approach.

### v2

@@ 5 Concluding remarks
In this paper the pure spinor model of type II superstring theory is studied. I choose to work with the superstring propagating in the presence of the constant background fields. The context of studying is T-dualization - combined bosonic and fermionic T-duality and double space representation of the full (bosonic++fermionic) T-dualization.
@@
At the beginning there are the separate recapitulations of the known results for bosonic and fermionic T-dualization of the type II superstring theory with constant background fields. The short discussion about choice of background fields is given and it is shown briefly the accordance with the consistency conditions. The main steps and results from [11] are repeated. Also in these papers [11] one can find more detailed analysis regarding the choice of the background fields.
@@
In the next part of the paper I showed that bosonic and fermionic Buscher T-dualization procedure, applied on the type II superstring with constant background fields, commute. Independent on sequence of the application of T-duality procedures, the result is unique - full T-dualized theory. Further, the background fields of the full T-dualized theory are obtained. The result is not trivial in the sense that the model analyzed in his paper is approximated - all background fields are constant. This result is an additional proof that such choice is correct.
@@
Final part of the paper is dedicated to the representation of full T-dualization in double space. Here I introduced the double coordinate and define the matrix of the full T-dualization, $\mathcal{T}^{M}{}_{N}$, which exchanges $x^{\mu}\leftrightarrow y_{\mu}$, $\theta^{\alpha}\leftrightarrow z_{\alpha}$ and $\bar{\theta}^{\alpha}\leftrightarrow\bar{z}_{\alpha}$, simultaneously. The full T-dualized coordinate is defined as $^{\star}Z^{M}=\mathcal{T}^{M}{}_{N}Z^{N}$. The next step was to rewrite the T-dual transformation laws using double coordinate and introducing so called generalized metric $\mathcal{H}_{\pm MN}$. Note that every chiral sector has its own generalized metric. Demanding that initial double coordinate $Z^{M}$ and fully T-dualized $^{\star}Z^{M}$ have the same T-dual transformation laws I get the form of the T-dualized generalized metric, $^{\star}\mathcal{H}_{\pm}=\mathcal{T}^{T}\mathcal{H}_{\pm}\mathcal{T}$. From one of these two equations I calculate the background fields of the fully T-dualized theory which are in full accordance with those obtained in the analytical (Buscher) procedure. For this specific choice of background fields I proved the equivalence of the analytical and double space approach.
At the end it is interesting to make comparison of the results from this paper with the recently published ones regarding double superspaces, double field theory and T-dualities. In the paper [17] it is proved that abelian Yang-Baxter (YB) deformations are equivalent to the commuting TsT transformations (T-dualities and coordinate shifts). Further, the result is generalised to fermionic deformations and fermionic T-duality, and finally to the super T-duality group $OSp(d_{b},d_{b}|2d_{f})$. This is in full correspondence with the result presented in this paper. On the other hand, Refs.[18, 19] considers the non-Abelian T-duality and its supersymmetric generalization. It is concluded taht there is a problem of integrating out all the initial variables in general case but in the case of the principal $OSp(1,2)$ model this problem disappears - super T-dualities commutes as it is the conclusion of this paper. In Refs.[20, 21, 22] it is considered the double field theory with the $OSp(d_{b},d_{b}|2d_{f})$ symmetry. As it is proved in [17], super YB deformations are equivalent to the commuting super T-dualizations.